using the formula s=2sqrt5x (I am having trouble inserting the radical sign so i used "sqrt" i know that is not right everything i used would not work). We will call the speed S (in miles per hour) and x is equal to the length of a skid mark in feet.
What is the speed of a car that has left a skid mark with a length of 61.25ft?

What is the speed of a car that has left a skid mark with a length of 320ft?

How long is the skid mark left by a car that was traveling 70 mph when involved in an accident?

Answer by Theo(13342)

your equation looks like it should be:
s = 2*sqrt(5x)
s = the speed of the car in miles per hour.
x = the length of the skid in feet.
when x = 61.5 feet, then:"
s = 2*sqrt(5*61.5) = 35.071 miles per hour
when x = 320 feet:
s = 2*sqrt(5*320) = 80 miles per hour.
a car traveling at 70 miles per hour would leave a skid mark as described in the following equation:
70 = 2*sqrt(5x)
divide this equation by 2 to get:
35 = sqrt(5x)
square both sides of this equation to get:
1225 = 5x
divide both sides of this equation by 5 to get:
x = 245 feet.
a car traveling at 70 miles per hour would leave a skid mark of 245 feet.
this assumes that it would skid to a stop without hitting anything.
if it hit something before coming to a stop, then the skid marks would be less than 245 feet.